1+2logx
Express as the log of a single quantity.

Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

1+2logx
Express as the log of a single quantity.
----------
Use parentheses.
1+2log(x)
= 1 + log(x^2)
= log(10) + log(x^2)
= log(10x^2)

Answer by greenestamps(13200)



No base is specified, so it is assumed the base is 10. Then

1 = log_10(10)
2log_10(x) = log_10(x^2)

Then
1+2logx
= 1+2log_10(x)
= log_10(10)+log_10(x^2)
= log_10[(10)*(x^2)]
= log_10(10x^2)
= log(10x^2)

Answer: 1+2logx = log(10x^2)